Gladow Corporation began operations in 2020 and has these transactions related to intangible assets in that year:
Jan. 2 Purchased a patent (5-year life) $276,000
Apr. 1 Goodwill purchased as part of an acquisition of another company $331,000
Jul. 1 Acquired a 9-year franchise; expiration date July 1, 2025 $496,800
Dec. 31 Determined that the recoverable amount of the patent and franchise be $161,000 and $483,000 respectively. There was no indication that the goodwill was impaired.

Collapse question part

(a)

Correct answer. Your answer is correct.
Prepare the necessary entries to record these intangibles. Assume all costs incurred were for cash. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Date

Account Titles and Explanation

Debit

Credit

Jan. 2

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Apr. 1

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

(Part of the entry to record
purchase of another company)

Jul. 1

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

Entry field with correct answer

SHOW LIST OF ACCOUNTS

LINK TO TEXT

LINK TO TEXT

LINK TO TEXT

Attempts: 1 of 2 used

Collapse question part

(b)

Prepare any adjusting journal entries required at December 31, 2020. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Date

Account Titles and Explanation

Debit

Credit

Dec. 31

(Amortization of Patent)

Dec. 31

(Amortization of Franchise)

Dec. 31

(Impairment of Patent)

Dec. 31

(Impairment of Franchises)

Dec. 31

(Impairment of Goodwill)

Part (b) only
Answer #1

Patent amortization per year=276000/5=55200
At end of year book value=276000-55200=$220,800
Recoverable amount of patent=161000
Impairement loss=220800-161000=59800

franchise rights amortization per month=496800/(9*12)=4600
Amortization for year=4600*6=27600
At end of year book value=496800-(4600*6)=$469,200
Recoverable amount of patent=483000
There is no impairment here

Date Account tilte and explanation Debit Credit Jan.2 Patent 276,000 Cash 276,000 Record the patent purchased for cash Apr.1